The Sales force for a business consists of 10 men and 10 women. A production unit of 7 people set up at random. what is the probability that it will consist of 4 men and 3 women.
Thank you

Answer by ewatrrr(24785)

 
Hi,
The Sales force for a business consists of 10 men and 10 women: 20 in ALL
A production unit of 7 people set up at random
P(4men and 3 women) = 10C4*10C3/20C7 = 210*120/77,520 = .3251